OpenStudy (anonymous):

Which statement is true? A)In any right triangle, the sine of one acute angle is equal to the cosine of the other acute angle B)In any right triangle, the sine of one acute angle is equal to the sine of its complementary angle. C)In any right triangle, the cosine of one acute angle is equal to the cosine of its complementary angle D)In any right triangle, the sum of the sine of one acute angle and the cosine of the other acute angle is 1 @chrisdbest

OpenStudy (chrisdbest):

That looks like B

OpenStudy (anonymous):

are you sure

OpenStudy (chrisdbest):

yes

OpenStudy (anonymous):

thanks

OpenStudy (chrisdbest):

medal please?

OpenStudy (anonymous):

it was wrong

OpenStudy (danjs):

Think about the values for X and Y in the unit circle, where (x,y) = (cos(theta), sin(theta))

OpenStudy (danjs):

For B as suggested, Test and see if \[\sin(\theta) = \sin(90- \theta)\]? are those two Y values the same if you use an angle less than 90 for theta...